# Automatic Fused Multimodal Deep Learning for Plant Identification

### v1

@@ Abstract
Plant classification is vital for ecological conservation and agricultural productivity, enhancing our understanding of plant growth dynamics and aiding species preservation. The advent of deep learning (DL) techniques has revolutionized this field by enabling autonomous feature extraction, significantly reducing the dependence on manual expertise. However, conventional DL models often rely solely on single data sources, failing to capture the full biological diversity of plant species comprehensively. Recent research has turned to multimodal learning to overcome this limitation by integrating multiple data types, which enriches the representation of plant characteristics. This shift introduces the challenge of determining the optimal point for modality fusion. In this paper, we introduce a pioneering multimodal DL-based approach for plant classification with automatic modality fusion. Utilizing the multimodal fusion architecture search, our method integrates images from multiple plant organs—flowers, leaves, fruits, and stems—into a cohesive model. Our method achieves 83.48% accuracy on 956 classes of the PlantCLEF2015 dataset, surpassing state-of-the-art methods. It outperforms late fusion by 11.07% and is more robust to missing modalities. We validate our model against established benchmarks using standard performance metrics and McNemar’s test, further underscoring its superiority.
@@ 1 Introduction
Contribution: In this paper, we propose a novel automatic fused multimodal DL approach for the first time in the context of plant classification. In this regard, the study first identifies the inadequacy of unimodal DL for plant classification, prompting an exploration of multimodal learning, particularly through multimodal fusion. Then, considering images of four visible plant organs – flowers, leaves, fruits, and stems as four different modalities, the automatic fusion of these modalities is performed using the multimodal fusion architecture search algorithm (MFAS) (Perez-Rua et al., 2019). To do so, an unimodal model for each modality is first trained using the MobileNetV3Small pre-trained model. These models are then fused using the MFAS algorithm, resulting in a unified model architecture. Moreover, we evaluate the proposed model against an established baseline in the form of late fusion using an averaging strategy (Baltrusaitis et al., 2019), utilizing standard performance metrics and McNemar’s statistical test (Dietterich, 1998). The results demonstrate that our automated fusion approach enables the construction of a more effective multimodal DL model for plant identification, outperforming the baseline. Additionally, we test our proposed model on individual plant organs, revealing that while expectedly unimodal models, specifically tackled for a single modality, perform better, the performance of the multimodal model is reasonably close to theirs. Finally, we assess the proposed model on subsets of plant organs to show its robustness to missing modalities. The source code of our work is available online at the GitHub repository [^1] GitHub repository: https://github.com/AlfredsLapkovskis/MultimodalPlantClassifier.
@@ 2.1 Fusion Automation Algorithm Selection
Different layers of DL models represent different levels of abstractions, and the highest levels are not necessarily the most suitable for fusion (Perez-Rua et al., 2019). Following this insight, to automate the development of a multimodal model architecture, we explore the use of NAS algorithms specifically designed for multimodal frameworks. These algorithms offer a promising solution for automatically identifying the optimal point of fusion rather than relying on manual determination. In this regard, Perez-Rua et al. (2019) introduce MFAS, a multimodal NAS algorithm. Central to their approach is the assumption that each modality possesses a distinct pre-trained model, substantially reducing the search space by maintaining these models static during the search process. The MFAS algorithm iteratively seeks an optimal joint architecture by progressively merging separate pre-trained models at layers. A notable advantage of this methodology lies in its focus on training fusion layers exclusively, resulting in significant computational time savings.
In another work, Xu et al. (2021) present MUFASA, an advanced multimodal NAS algorithm. One of the standout features of this algorithm is its comprehensive approach, which involves searching for optimal architectures not only for the entire fusion architecture but also for each modality individually, all while considering various fusion strategies. Unlike unimodal NAS methods, MUFASA addresses the whole architecture, leveraging the understanding of multimodal nature. Furthermore, unlike Perez-Rua et al.’s (2019) algorithm, MUFASA addresses the architectures of individual modalities while considering their interdependencies. This unique approach positions MUFASA as potentially more powerful at tackling challenges in multimodal fusion.
@@ 2.2.1 Unimodal Models
To construct a unimodal model for each modality represented by different plant organs, we employ a transfer learning technique. There are two primary approaches to applying transfer learning. One involves utilizing pre-trained model weights to extract features from a dataset for subsequent classification, while the other entails full or partial updates of pre-trained weights using a new dataset, known as fine-tuning (Espejo-Garcia et al., 2020). We rely on the former approach, leveraging pre-trained weights and replacing a classification layer, as we did not observe an improvement from fine-tuning. Thus, only the top layer undergoes training.
@@ 2.2.2 Multimodal Fusion Architecture Search
We leverage MFAS algorithm, which utilizes a pre-trained model $\bm{f}^{(i)}(\bm{x}^{(i)})=\hat{z_{i}}$ for each modality $i\in\{1,\dots,m\}$. Here, $\hat{z_{i}}$ denotes an approximation of the true label $z$, derived from the input $\bm{x}$ specific to modality $i$. Each model $\bm{f}^{(i)}$ with size $n_{i}$ is composed of layers $\bm{f}^{(i)}_{l}$, where $l\in\{1,\dots,n_{i}\}$, such that for $l$-th layer, $\bm{x}^{(i)}_{l}=(\bm{f}^{(i)}_{l}\circ\bm{f}^{(i)}_{l-1}\circ\dots\circ\bm{f}^{(i)}_{1})(\bm{x}^{(i)})$ represents features considered for fusion with features from other modalities. The objective of the algorithm is to find optimal combinations of such features to fuse, along with determining the properties of such fusion.
@@
$\bm{h}_{l}=\begin{cases}\bm{\sigma}_{\gamma^{a}_{1}}\begin{pmatrix}\bm{W}_{1}\begin{bmatrix}\bm{x}^{(1)}_{\gamma^{1}_{1}}\\ \vdots\\ \bm{x}^{(m)}_{\gamma^{m}_{1}}\\ \end{bmatrix}\end{pmatrix}&\text{if }l=1\\ \\ \bm{\sigma}_{\gamma^{a}_{l}}\begin{pmatrix}\bm{W}_{l}\begin{bmatrix}\bm{x}^{(1)}_{\gamma^{1}_{l}}\\ \vdots\\ \bm{x}^{(m)}_{\gamma^{m}_{l}}\\ \bm{h}_{l-1}\\ \end{bmatrix}\end{pmatrix}&\text{if }l>1\end{cases}$ (1)
where for $l$-th layer of $\bm{h}$, $\bm{\sigma}_{\gamma^{a}_{l}}$ denotes an activation function, $\bm{W}_{l}$ is a trainable weight matrix and $\gamma_{l}=(\gamma^{1}_{l},\dots,\gamma^{m}_{l},\gamma^{a}_{l})$ is a tuple with indices of features from each modality and an index of an activation function. The maximum number of fusion layers $\bm{h}_{l}$ is denoted by $L:l\in\{1,\dots,L\}$, and is a hyperparameter defined prior to execution of the algorithm. A complete fusion configuration of a particular instance of $\bm{h}$ with $L$ layers is defined by a vector of tuples $[\gamma_{l}]_{l\in\{1,...,L\}}$, while a set of all possible tuples with $L$ layers is denoted by $\bm{\Gamma}_{L}$. A list of possible activation functions with size $k$ and possible modality layers used in $\bm{h}_{l}$ are also hyperparameters that can be implementation-specific.
@@
  • Unimodal models. We utilize the pre-trained MobileNetV3Small model for each modality as detailed in Section 2.2.1.
  • Search space. In this paper, the search space is constrained to the following parameters:
    • Maximum of three fusion layers (i.e., $L=3$).
@@
    • 5 fusible layers in each unimodal model: the 1st activation, the 1st, 6th, and 11th inverted residual blocks, and the output from the global average pooling layer on the top of the model (i.e., $n_{i}=5$).
    This results in $12505^{4}\times 2=$ 1250 different configurations for a single fusion layer, and forms a search space of size $1250^{3}\approx 1.95\times 10^{9}$.
  • Generating model configurations. In this work, the sampled architectures, denoted by $\mathcal{S}$, are reused across iterations, in contrast to the approach taken by Perez-Rua et al. (2019), where architectures are regenerated in each iteration. During the progression from $l=1$ to $l=L$, newly generated layer configurations are either appended to the sampled configurations or used to replace existing $l$-th layers in $\mathcal{S}$.
  • Building architectures. Multimodal architectures are constructed based on fusion configurations. Since the outputs of layers from unimodal models may have different numbers of dimensions, similar to Perez-Rua et al. (2019), global average pooling is applied to each multidimensional output from MobileNetV3Small. Subsequently, the output of each fused layer is concatenated and connected with a dense layer, incorporating an activation function specified in the model configuration. In the case of an intermediate fusion layer ($l>1$), the previous fusion layer is also connected to the dense layer. Each multimodal model is finalized with a classifier layer employing a softmax activation function.
  • Weight sharing. In contrast to Perez-Rua et al. (2019), where weights are shared between fusion layers that have the same indices and identical weight matrix sizes, our approach also takes into account the activation functions of the fusion layers. This adjustment acknowledges the significant impact that activation functions can have on the behavior of weights. In our model, weights are shared across all configurations.
  • Storing results. Similar to Perez-Rua et al. (2019), if the same configuration is visited twice, the best result is retained.
  • Surrogate. In our work, a surrogate similar to the one employed by Perez-Rua et al. (2018) is utilized, as it has proven to be effective in this context. This surrogate comprises an embedding layer with zero masking, yielding vectors with a length of 100; an LSTM layer with 100 neurons; and a regression layer with a single neuron and sigmoid activation. The model is compiled with an Adam optimizer with a learning rate of 0.001, and mean squared error (MSE) loss. Each update is executed for 50 epochs with a batch size of 64.
@@
  • Temperature-based sampling. According to Perez-Rua et al. (2018), the probability of sampling an architecture $i$ with a score $a_{i}$ is $p_{i}=a_{i}/\sum_{j}a_{j}$. However, when incorporating a temperature factor $t$, the probability of sampling $i$ is given by $p_{i}^{1/t}/\sum_{j}p_{j}^{1/t}$. Thus, the larger the $t$, the more stochastic the sampling becomes. Similar to Perez-Rua et al. (2019), this research employs inverse exponential temperature scheduling, as it has shown to perform well in this context too. Therefore, at each step $s$ of the MFAS algorithm, the temperature is computed as $(t_{max}-t_{min})e^{-(s/d)^{2}}+t_{min}$.
@@ 2.3 Dataset
We selected the PlantCLEF2015 dataset (Joly et al., 2015) for our proposed plant classification model because it includes images of different plant organs— including leaves, flowers, fruits, and stems—providing a large and diverse collection that ensures a sufficient quantity of images for each organ type as depicted in Figure 1.
Figure 1: Plant Organ Distributions in PlantCLEF2015
This dataset follows a long-tailed distribution, as depicted in Figure 2. This class imbalance in the data underscores the necessity of employing additional techniques to mitigate its potentially adverse effects on model learning. Furthermore, Figure 3 illustrates the distribution of organ images across classes. Notably, a considerable number of classes entirely lack certain organs. This observation highlights the necessity for models to handle missing modalities effectively, necessitating tailored training approaches.
Figure 2: Class Distribution in PlantCLEF2015
Only the relevant organs—leaves, flowers, fruits, and stems—are considered.
Figure 3: Organ Distribution among Classes in PlantCLEF2015
Logarithmic scale is used to visualize low values.
Moreover, it is evident that flowers, being a discriminative organ (Nhan et al., 2020), are available for most classes, followed by leaves and fruits, although with a noticeable scarcity in many classes. This scarcity is even more pronounced in stems, being the least discriminative and the least available organ across classes. These insights emphasize the importance of addressing the distributional characteristics of organ images within the dataset for robust model training and classification performance.
@@ 2.3.1 Data Preprocessing
    Collecting metadata. Initially, we extract image metadata from the dataset.
@@
    Standardize images. We convert each image into RGB representation and resize it to the $256\times 256$ resolution. This standardization ensures uniform model input dimensions and minimizes computational load. The adoption of a square shape for resizing is justified by the observation that the average aspect ratio of images in the dataset closely approximates 1.
@@
    Grouping image indices and corresponding labels. We iterate the metadata to group image indices and labels based on the organs they depict. This is a crucial step for the subsequent modality-wise preprocessing.
@@
    Filtering small classes. A number of classes in the dataset contain too low a number of images for effective classification, therefore, in this step, we exclude small classes. The classes are filtered in each modality separately because each modality is used individually in step 6, and the proposed multimodal model uses all four modalities simultaneously. A threshold of a minimum of 10 organ images per class is established to include a class in a modality. This filtering results in 302 classes in fruit, 449 in leaf, 862 in flower and 146 in stem modalities, and 956 unique classes in total.
@@
    Splitting organ indices and corresponding labels. To facilitate model training and evaluation, the data undergoes a stratified split into training, validation, and test sets, with proportions of $60\%:20\%:20\%$. Maintaining consistency with the earlier approaches, splitting is conducted at the modality level.
@@
    Creating references to unimodal data. Following the observation splitting in step 5, pairs of image indices and labels are organized into data frames to enable loading images and corresponding labels during training, validation and testing of unimodal models (Section 2.2.1).

### v2

@@ Abstract
Plant classification is vital for ecological conservation and agricultural productivity, enhancing our understanding of plant growth dynamics and aiding species preservation. The advent of deep learning (DL) techniques has revolutionized this field by enabling autonomous feature extraction, significantly reducing the dependence on manual expertise. However, conventional DL models often rely solely on single data sources, failing to capture the full biological diversity of plant species comprehensively. Recent research has turned to multimodal learning to overcome this limitation by integrating multiple data types, which enriches the representation of plant characteristics. This shift introduces the challenge of determining the optimal point for modality fusion. In this paper, we introduce a pioneering multimodal DL-based approach for plant classification with automatic modality fusion. Utilizing the multimodal fusion architecture search, our method integrates images from multiple plant organs—flowers, leaves, fruits, and stems—into a cohesive model. Our method achieves 82.61% accuracy on 979 classes of the PlantCLEF2015 dataset, surpassing state-of-the-art methods and outperforming late fusion by 10.33%. Through the incorporation of multimodal dropout, our approach demonstrates strong robustness to missing modalities. We validate our model against established benchmarks using standard performance metrics and McNemar’s test, further underscoring its superiority.
@@ 1 Introduction
Contribution: In this paper, we propose a novel automatic fused multimodal DL approach for the first time in the context of plant classification. In this regard, the study first identifies the inadequacy of unimodal DL for plant classification, prompting an exploration of multimodal learning, particularly through multimodal fusion. Then, considering images of four visible plant organs—flowers, leaves, fruits, and stems—as distinct modalities, we apply the multimodal fusion architecture search algorithm (MFAS) (Perez-Rua et al., 2019) to automatically fuse these modalities. To do so, a unimodal model for each modality is first trained using the MobileNetV3Small pre-trained model. These models are then fused using the MFAS algorithm, resulting in a unified model architecture. Moreover, we evaluate the proposed model against an established baseline in the form of late fusion using an averaging strategy (Baltrusaitis et al., 2019), utilizing standard performance metrics and McNemar’s statistical test (Dietterich, 1998). The results demonstrate that our automated fusion approach enables the construction of a more effective multimodal DL model for plant identification, outperforming the baseline. Additionally, we assess the proposed model on all subsets of plant organs, revealing its robustness to missing modalities when trained with a multimodal dropout technique (Cheerla and Gevaert, 2019). The source code of our work is available online at the GitHub repository [^1] GitHub repository: https://github.com/AlfredsLapkovskis/MultimodalPlantClassifier.
@@ 2.1 Fusion Automation Algorithm Selection
Different layers of DL models represent different levels of abstractions, and the highest levels are not necessarily the most suitable for fusion (Perez-Rua et al., 2019). Following this insight, to automate the development of a multimodal model architecture, we explore the use of NAS algorithms specifically designed for multimodal frameworks. These algorithms offer a promising solution for automatically identifying the optimal point of fusion rather than relying on manual determination. In this regard, Perez-Rua et al. (2019) introduce MFAS, a multimodal NAS algorithm. Central to their approach is the assumption that each modality possesses a distinct pre-trained model, substantially reducing the search space by maintaining these models static during the search process. The MFAS algorithm iteratively seeks an optimal joint architecture by progressively merging individual pre-trained models at different layers. A notable advantage of this methodology lies in its focus on training fusion layers exclusively, resulting in significant computational time savings.
In another work, Xu et al. (2021) present MUFASA, an advanced multimodal NAS algorithm. One of the standout features of this algorithm is its comprehensive approach, which involves searching for optimal architectures not only for the entire fusion architecture but also for each modality individually, all while considering various fusion strategies. Unlike unimodal NAS methods, MUFASA addresses the whole architecture, leveraging the understanding of its multimodal nature. Furthermore, unlike Perez-Rua et al.’s (2019) algorithm, MUFASA addresses the architectures of individual modalities while considering their interdependencies. This unique approach positions MUFASA as potentially more powerful at tackling challenges in multimodal fusion.
@@ 2.2.1 Unimodal Models
To construct a unimodal model for each modality represented by different plant organs, we employ a transfer learning technique. There are two primary approaches to applying transfer learning. One involves utilizing pre-trained model weights to extract features from a dataset for subsequent classification, while the other entails full or partial updates of pre-trained weights using a new dataset, known as fine-tuning (Espejo-Garcia et al., 2020). Initially, we adopted the former approach, pre-training only the appended top layers of each unimodal model. For models where fine-tuning yielded improved performance, we selectively proceeded with this technique.
@@ 2.2.2 Multimodal Fusion Architecture Search
We leverage the MFAS algorithm, which utilizes a pre-trained model $\bm{f}^{(i)}(\bm{x}^{(i)})=\hat{y_{i}}$ for each modality $i\in\{1,\dots,m\}$. Here, $\hat{y_{i}}$ denotes an approximation of the true label $y$, derived from the input $\bm{x}$ specific to modality $i$. Each model $\bm{f}^{(i)}$ with size $n_{i}$ is composed of layers $\bm{f}^{(i)}_{l}$, where $l\in\{1,\dots,n_{i}\}$, such that for $l$-th layer, $\bm{x}^{(i)}_{l}=(\bm{f}^{(i)}_{l}\circ\bm{f}^{(i)}_{l-1}\circ\dots\circ\bm{f}^{(i)}_{1})(\bm{x}^{(i)})$ represents features considered for fusion with features from other modalities. The objective of the algorithm is to find optimal combinations of such features to fuse, along with determining the properties of such fusion.
@@
$\bm{h}_{1}=\bm{\sigma}_{\gamma^{a}_{1}}\begin{pmatrix}\bm{W}_{1}\begin{bmatrix}\bm{x}^{(1)}_{\gamma^{1}_{1}}\\ \vdots\\ \bm{x}^{(m)}_{\gamma^{m}_{1}}\\ \end{bmatrix}\end{pmatrix},\bm{h}_{l}=\bm{\sigma}_{\gamma^{a}_{l}}\begin{pmatrix}\bm{W}_{l}\begin{bmatrix}\bm{x}^{(1)}_{\gamma^{1}_{l}}\\ \vdots\\ \bm{x}^{(m)}_{\gamma^{m}_{l}}\\ \bm{h}_{l-1}\\ \end{bmatrix}\end{pmatrix},l>1$ (1)
where for $l$-th layer of $\bm{h}$, $\bm{\sigma}_{\gamma^{a}_{l}}$ denotes an activation function, $\bm{W}_{l}$ is a trainable weight matrix and $\gamma_{l}=(\gamma^{1}_{l},\dots,\gamma^{m}_{l},\gamma^{a}_{l})$ is a tuple with indices of features from each modality and an index of an activation function. The maximum number of fusion layers $\bm{h}_{l}$ is denoted by $L:l\in\{1,\dots,L\}$, and is a hyperparameter defined prior to execution of the algorithm. A complete fusion configuration of a particular instance of $\bm{h}$ with $L$ layers is defined by a vector of tuples $[\gamma_{l}]_{l\in\{1,\dots,L\}}$, while a set of all possible tuples with $L$ layers is denoted by $\bm{\Gamma}_{L}$. A list of possible activation functions with size $k$ and possible modality layers used in $\bm{h}_{l}$ are also hyperparameters that can be implementation-specific.
@@
  • Unimodal models: We utilize the pre-trained MobileNetV3Small model for each modality as detailed in Section 2.2.1.
  • Search space: In this paper, the search space is constrained to the following parameters:
    • Maximum of four fusion layers (i.e., $L=4$).
@@
    • 6 fusible layers in each unimodal model: the 1st activation, the 1st, 6th, and 11th inverted residual blocks, and our two appended dense layers (i.e., $n_{i}=6,\forall i$).
    This results in $25926^{4}\times 2=$ 2592 different configurations for a single fusion layer, and forms a search space of size $2592^{4}\approx 4.51\times 10^{13}$.
  • Generating model configurations: In this work, the sampled architectures, denoted by $\mathcal{S}$, are reused across iterations, in contrast to the approach taken by Perez-Rua et al. (2019), where architectures are regenerated in each iteration. During the progression from $l=1$ to $l=L$, newly generated layer configurations are either appended to the sampled configurations or used to replace existing $l$-th layers in $\mathcal{S}$.
  • Building architectures: Multimodal architectures are constructed based on fusion configurations. Since the outputs of layers from unimodal models may have different numbers of dimensions, similar to Perez-Rua et al. (2019), global average pooling is applied to each multidimensional output from MobileNetV3Small. Subsequently, the output of each fused layer is concatenated and connected with a dense layer, incorporating an activation function specified in the model configuration. In the case of an intermediate fusion layer ($l>1$), the previous fusion layer is also connected to the dense layer. Each multimodal model is finalized with a classifier layer employing a softmax activation function.
  • Weight sharing: In contrast to Perez-Rua et al. (2019), where weights are shared between fusion layers that have the same indices and identical weight matrix sizes, our approach also takes into account the activation functions of the fusion layers. This adjustment acknowledges the significant impact that activation functions can have on the behavior of weights. In our model, weights are shared across all configurations.
  • Storing results: Similar to Perez-Rua et al. (2019), if the same configuration is visited twice, the best result is retained.
  • Surrogate: In our work, a surrogate similar to the one employed by Perez-Rua et al. (2018) is utilized, as it has proven to be effective in this context. This surrogate comprises an embedding layer with zero masking, yielding vectors with a length of 100; an LSTM layer with 100 neurons; and a regression layer with a single neuron and sigmoid activation. The model is compiled with an Adam optimizer with a learning rate (LR) of 0.001, and mean squared error (MSE) loss. Each update is executed for 50 epochs with a batch size of 64.
@@
  • Temperature-based sampling: According to Perez-Rua et al. (2018), the probability of sampling an architecture $i$ with a score $a_{i}$ is $p_{i}=a_{i}/\sum_{j}a_{j}$. However, when incorporating a temperature factor $t$, the probability of sampling $i$ is given by $p_{i}^{1/t}/\sum_{j}p_{j}^{1/t}$. Thus, the larger the $t$, the more stochastic the sampling becomes. Similar to Perez-Rua et al. (2019), this research employs inverse exponential temperature scheduling, as it has shown to perform well in this context too. Therefore, at each step $s$ of the MFAS algorithm, the temperature is computed as $(t_{max}-t_{min})e^{-(s/d)^{2}}+t_{min}$, where $d$ is the decay rate.
@@ 2.3 Dataset
We selected the PlantCLEF2015 dataset (Joly et al., 2015) for our proposed plant classification model because it contains images of different plant organs—including flowers, leaves, fruits, and stems—providing a large and diverse collection that ensures a sufficient quantity of images for each organ type as depicted in Figure 1(a).
Figure 1: PlantCLEF2015 Dataset Summary
(a) Distribution of plant organs in the dataset, (b) Class distribution in the dataset, considering only the relevant organs: flowers, leaves, fruits, and stems.
This dataset follows a long-tailed distribution, as depicted in Figure 1(b). This class imbalance in the data underscores the necessity of employing additional techniques to mitigate its potentially adverse effects on model learning. Furthermore, Figure 2 illustrates the distribution of organ images across classes. Notably, a considerable number of classes entirely lack certain organs. This observation highlights the necessity for models to handle missing modalities effectively, necessitating tailored training approaches.
Figure 2: Distribution of Plant Organs among Classes in PlantCLEF2015
Moreover, it is evident that flowers, being a discriminative organ (Nhan et al., 2020), are available for most classes, followed by leaves and fruits, although with a noticeable scarcity in many classes. This scarcity is even more pronounced in stems, being the least discriminative and the least available organ across classes.
Finally, the images in the dataset are derived from observations—individual plant specimens. It is essential to use this information to split the data for training, validation, and evaluation based on observations rather than images, in order to prevent potential bias from exposing similar images of the same observation across different data splits. However, achieving well-balanced splits in this context is non-trivial, as each observation can have an arbitrary number of images, as illustrated in Figure 3. Our approach to data splitting under these circumstances is detailed in Section 2.3.1.
Figure 3: Distribution of Plant Organs among Observations in PlantCLEF2015
@@ 2.3.1 Data Preprocessing
    Combining the original splits: The original data splits of PlantCLEF2015 are poorly balanced with respect to plant organs; therefore, we merge them into a unified dataset to generate our own splits subsequently.
@@
    Collecting metadata: Initially, we extract image metadata for flowers, leaves, fruits and stems from the dataset, resulting in a total of 67812 records.
@@
    Grouping class observations: We group the metadata of different organ images by their observation identifiers for each class separately.
@@
    Filtering organ images and classes: We eliminate specific organ images from a class if their total count in this class is less than 3, as such images cannot be split between training, validation and test sets. Subsequently, we remove any observations that are left empty. Finally, we discard classes with fewer than 3 observations, since we split the data by observations (Section 2.3). This filtering results in 919 classes for flowers, 726 for leaves, 500 for fruits, and 314 for stems, yielding a total of 979 unique classes.
@@
    Splitting the data: We split the data by observations; however, achieving balanced splits is challenging as (i) we aim to ensure a specific proporion of observations in each data split—i.e., 60%:20%:20% for training, validation and test sets respectively—to maximize the diversity of data within each split, and, simultaneously, (ii) we seek to ensure the same proportions of each plant organ’s images in the splits to achieve certain split sizes and balance in terms of organs. These objectives are inherently contradictory, as observations contain varying numbers of organ images (Fig. 3). Consequently, we frame the splitting process as a constrained optimization problem, which we solve for each class separately to ensure stratification by labels:
    minimize $\begin{split}\textrm{minimize\hskip 10.00002pt}&\underbrace{\sum_{s\in S}(||\mathbf{x}_{s}||_{1}-\lambda_{s}N)^{2}}_{\text{optimize observations}}+\underbrace{\sum_{o\in O}\sum_{s\in S}(\mathbf{c}_{o}^{T}\mathbf{x}_{s}-\lambda_{s}||\mathbf{c}_{o}||_{1})^{2}}_{\text{optimize organ counts}}\\ \textrm{subject to\hskip 10.00002pt}&\sum_{s\in S}\mathbf{x}_{s}=\mathbf{1}\end{split}$ (2)
    where $N$ is the number of observations, $S$ is a set of data splits, $O$ is the set of organ types, $\mathbf{c}_{o}\in\mathbb{N}^{N}$ is a vector containing the counts of organ $o$ images for each observation, $\mathbf{x}_{s}\in\{0,1\}^{N}$ is a decision vector for a split $s$, where the $i$-th element is 1 if the $i$-th observation belongs to split $s$, and $\lambda_{s}\in[0,1]$ is the desired size of split $s$ ($0.6$ for training and $0.2$ for validation and test).
    After solving this problem, we assign observations to splits according to the resulting vectors $\mathbf{x}_{s}$. If an organ split in a particular class ends up empty because all images were assigned to other splits, we manually transfer a single image from the largest split to ensure that training, validation, and testing can be performed for that class. This correction is practically expectable when a class has, for example, 3–4 images of the organ. Overall, this procedure produces reasonably balanced splits, as shown in Figure 4. Table 1 summarizes the total counts of observations and organ images across the splits.
@@
    Saving unimodal datasets: We save each data split separately for each modality to enable pre-training of unimodal models for MFAS. Each dataset consists of shuffled images and their corresponding labels, with labels mapped to values in the range $[0,978]$, as the original label values are impractical for use. Each image is converted to RGB, resized to $256\times 256$ resolution and encoded in JPEG format. This standardization ensures uniform model input dimensions and reduces computational and I/O load. The choice of a square shape for resizing is justified by the observation that the average aspect ratio of images in the dataset closely approximates 1. During data loading, we also normalize pixel channel values to the range $[-1,1]$ to meet the input requirements of MobileNetV3Small.
@@
Splits: Training Validation Test
Observations 16737 (59.21%) 5788 (20.48%) 5740 (20.31%)
Flowers 21859 (60.09%) 7273 (19.99%) 7243 (19.91%)
Leaves 9357 (59.84%) 3172 (20.29%) 3107 (19.87%)
Fruits 5096 (59.97%) 1678 (19.75%) 1724 (20.29%)
Stems 3196 (59.05%) 1119 (20.68%) 1097 (20.27%)
Table 1: Counts of Observations and Organs in Data Splits
Figure 4: Distribution of Organs across Different Splits